Consider a step-index fiber for which n1 = 1.45, n2 = 1.24 and a = 24µm. What is the maximum value of numerical aperture.


1
Expert's answer
2021-05-09T13:06:57-0400

"\sin \theta=\sqrt{n_1^2-n_2^2},"

"\sin\theta=\sqrt{1.45^2-1.24^2}=0.752,"

"\theta=48.7°."